Can I get some help on these problems please. Thank you very much.
Find the slope of the line passing through the points(-8,-3) and (-2,2).
Which of the ordered pairs
(3,1),(0,-4),(-4,0),(-3,-7) are solutions for the equation x-y=4?
An employee who produces x units per hour earns an hourly wage of y=0.35x+11(in dollars). What is the hourly wage of an employee who produces 2 units per hour?
graph x=3.

Thank you for your time.

Answer by vertciel(183)







New Page 1


1.
Find the slope of the line passing through the points(-8,-3) and (-2,2).


 


Slope =
(Y2 - Y1)/(X2 - X1)


Just
identify which coordinate corresponds to which in the slope formula, substitute
them and you will have your slope.




2. Which of the ordered pairs

(3,1),(0,-4),(-4,0),(-3,-7) are solutions for the equation x-y=4?


You can
rewrite x -y = 4 as


x = 4 +
y


or


x - 4 =
y


I'll go
with x = 4 + y since it's easier to use. This equation states the value of 4 and
the value of the y coordinate gives x. So we have to use guess and check here,
but it's really quick to do. (3,1): Does 4 + 1 = 3?  (0,-4): Does 4 + -4 =
0? Continue with this logic and you will have your answer(s). (There may be more
than one answer.)




3. An employee who produces x units per hour earns an hourly wage of
y=0.35x+11(in dollars). What is the hourly wage of an employee who produces 2
units per hour?



X is the units per hour, and since you want to find out the hourly wage of an
employee who produces 2 units per hour, just substitute 2 into the formula:


y =
0.35(2) + 11


Y
represents dollar so just solve for y.


 


4.
Graph x = 3


You
know that if x = 3, this line must run through the point (3,0). Now all you need
is the slope. However, is there slope here? If there isn't, what does that
mean?